OpenStudy (anonymous):

help me solve pleeease ! log x=4

OpenStudy (anonymous):

x = 10^4

OpenStudy (anonymous):

if natural log then e^4

OpenStudy (anonymous):

so 10,000? thank you!

OpenStudy (anonymous):

log(x) in general means log of x to the base 10, unless a base has been specifically mentioned. so, here, x = 10^4 = 10,000